OpenStudy (anonymous):

need help with system: -14=-2=y-7x 2x=10y +4 I have 20y-4x=-8 and -20y-7x=-14, but i think i did something wrong after that

OpenStudy (anonymous):

Try solving for y or x in both equations first.

OpenStudy (anonymous):

i got -11x=-24

OpenStudy (anonymous):

Show me the original problem again, please?

OpenStudy (anonymous):

-14=-2=y-7x 2x=10y +4

OpenStudy (anonymous):

-14 = -2. . . ?

OpenStudy (anonymous):

sorry -14=-20y-7x

OpenStudy (anonymous):

well if solving for y, -14=-20y-7x Since y is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. -20y-7x=-14 Since -7x does not contain the variable to solve for, move it to the right-hand side of the equation by adding 7x to both sides. -20y=7x-14 Divide each term in the equation by -20. -(20y)/(-20)=(7x)/(-20)-(14)/(-20) Simplify the left-hand side of the equation by canceling the common factors. y=(7x)/(-20)-(14)/(-20) Simplify the right-hand side of the equation by simplifying each term. y=-(7x)/(20)+(7)/(10)

OpenStudy (anonymous):

but if solving for x, -14=-20y-7x Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. -20y-7x=-14 Since -20y does not contain the variable to solve for, move it to the right-hand side of the equation by adding 20y to both sides. -7x=20y-14 Divide each term in the equation by -7. -(7x)/(-7)=(20y)/(-7)-(14)/(-7) Simplify the left-hand side of the equation by canceling the common factors. x=(20y)/(-7)-(14)/(-7) Simplify the right-hand side of the equation by simplifying each term. x=-(20y)/(7)+2

OpenStudy (anonymous):

and on the second half of the problem u listed, solving for y, 2x=10y+4 Since y is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 10y+4=2x Since 4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4 from both sides. 10y=-4+2x Move all terms not containing y to the right-hand side of the equation. 10y=2x-4 Divide each term in the equation by 10. (10y)/(10)=(2x)/(10)-(4)/(10) Simplify the left-hand side of the equation by canceling the common factors. y=(2x)/(10)-(4)/(10) Simplify the right-hand side of the equation by simplifying each term. y=(x-2)/(5)

OpenStudy (anonymous):

but solving for x would give you, 2x=10y+4 Divide each term in the equation by 2. (2x)/(2)=(10y)/(2)+(4)/(2) Simplify the left-hand side of the equation by canceling the common factors. x=(10y)/(2)+(4)/(2) Simplify the right-hand side of the equation by simplifying each term. x=5y+2

OpenStudy (anonymous):

take your pick on how you want to do the answer's